OpenStudy (anonymous):

is the common denominator for 2/4 + 1/5 20 ?

OpenStudy (anonymous):

yes

OpenStudy (aroub):

yes!! :)

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

so the next equation should look like 10/20 because 4*5=20 2*5=10

OpenStudy (anonymous):

now 1*5=5 5*5=25 am i going in the right direction ?

OpenStudy (aroub):

yes in the right direction!

OpenStudy (anonymous):

example 3/4 + 1/6 =9/12+2/12 =11/12 where did the 2 come from ?

OpenStudy (aroub):

oh... because u multiplied the 2 by 1/6 to have a common denominator which is 12 right?

OpenStudy (anonymous):

yes

OpenStudy (aroub):

yes thats why then! :)

OpenStudy (aroub):

do u have any questions abt this?

OpenStudy (aroub):

im free to help! :)

OpenStudy (anonymous):

yea so like with 2/4+1/5 id use this same method

OpenStudy (aroub):

yes the exact same method!! :)

OpenStudy (aroub):

if u want solve it and ill check it for you! :)

OpenStudy (anonymous):

ok so far i have 2/4 +1/5 =10/20 2*5=10 4*5=20 now 1*5=5 5*5=25 now im stuck ?help